OpenStudy (anonymous):

The perimeter of a rectangle is 70 m the length is 8 m more than twice the width find the dimensions

OpenStudy (anonymous):

P = L + L + W + W = 2L + 2W P = 70m 70 = 2L + 2W L = (8 + 2W) So substitute, L inside the Perimeter equation 70 = 2(8+2W) + 2W Solve for W, then plug back into the Length equation to find L.

OpenStudy (anonymous):

W = 9m , L = 26m